# Torsion and Chern-Simons gravity in 4D space-times from a Geometrodynamical four-form

### v1

@@ Abstract
By envisaging a four dimensional manifold’s space-time geometry be described by a line-element, $ds^{2}=\eta_{\mu\nu}dx^{\mu}dx^{\nu}$, Hermann Minkowski had overturned the idea of treating space and time disjointly. Furthermore, the cause and effect relation, in the realm of classical physics, attained a significant elucidation in the Minkowskian framework - infinitesimally separated events (assuming a signature (+ - - -)) for which $ds^{2}<0$ ($ds^{2}>0$) are causally disconnected (causally connected). Of course, the underlying physics involved in it is that an effect essentially follows a cause due to transfer of energy from the latter to the former, and that the rate of flow of energy is limited by the speed $c$. What is usually not stressed is that, because of this causality criteria, the imaginary numbers entered physics in a fundamental manner even before the advent of quantum mechanics: the proper distance $ds$ between any two causally disconnected events is always an imaginary number.
@@
It must be pointed out that there is another tensor in the flat space-time that is invariant under proper Lorentz transformations - the totally antisymmetric Levi-Civita tensor $\epsilon_{\mu\nu\rho\sigma}$. Since, under $x^{\mu}\rightarrow\Lambda^{\mu}_{\ \nu}x^{\nu}$,
@@
and $det(\Lambda)=1$ for proper Lorentz Transformations, it is clear that the flat space-time Levi-Civita tensor does not change under arbitrary boosts and rotations of inertial frame coordinates.
Given this special status that the Minkowski metric and the Levi-Civita tensor enjoy in the Minkowskian space-time, as well as the fact that $\eta_{\mu\nu}$ metamorphoses into a dynamical space-time metric $g_{\mu\nu}(x)$ in GR, raises a pertinent point of of elevating $\epsilon_{\mu\nu\alpha\beta}$ into a dynamical field. In the present article, we explore such a possibility by positing a geometrodynamical field, $w_{\mu\nu\rho\sigma}(x)$, totally antisymmetric in its indices, that is independent of the metric tensor and that influences space-time physics so that an interesting status that $\eta_{\mu\nu}$ and $\epsilon_{\mu\nu\alpha\beta}$ both share in the flat space-time is true also in curved space-times pdg-ref-20 .
We propose in this article that this geometrodynamical field represents a new physical degree of freedom that couples universally to all matter. In the following section, we discuss some of the differential geometric aspects of $w_{\mu\nu\rho\sigma}(x)$ that may be significant in the context of space-time geometry.
@@
In a coordinate basis, the corresponding four-form field may be expressed as,
@@
Since we demand that this new dynamical field be independent of the metric tensor, instead of raising the indices of $w_{\mu\nu\rho\sigma}$ using the metric tensor, the totally antisymmetric components $w^{\mu\nu\rho\sigma}$ corresponding to the four-form $\tilde{w}$ are obtained from the associated p-vector $\mathcal{\omega}$, so that one obtains the relation, $w^{\mu\nu\rho\sigma}\ w_{\mu\nu\rho\sigma}=-4!$ Schutz80 . This implies,
@@
$\bar{\Gamma}^{\mu}_{\alpha\beta}\equiv\Gamma^{\mu}_{\alpha\beta}-a_{T}\ g^{\mu\nu}\chi^{\ ;\gamma}w_{\nu\gamma\alpha\beta}\ ,$ (10)
@@
Covariant derivatives of tensor fields are obtained in the standard manner by replacing $\Gamma^{\mu}_{\alpha\beta}$ with $\bar{\Gamma}^{\mu}_{\alpha\beta}$ of eq.(10) so that,
@@
$D_{\nu}B_{\mu}=\partial_{\nu}B_{\mu}-\bar{\Gamma}^{\beta}_{\mu\nu}B_{\beta}=B_{\mu;\ \nu}-g^{\beta\sigma}\chi^{;\gamma}w_{\sigma\gamma\mu\nu}B_{\beta}\ \ ,$ (13)
@@
$D_{\nu}C_{\mu\tau}=\partial_{\nu}C_{\mu\tau}-\bar{\Gamma}^{\beta}_{\mu\nu}C_{\beta\tau}-\bar{\Gamma}^{\beta}_{\tau\nu}C_{\mu\beta}=C_{\mu\tau;\ \nu}-g^{\beta\sigma}\chi^{;\gamma}w_{\sigma\gamma\mu\nu}C_{\beta\tau}-g^{\beta\sigma}\chi^{;\gamma}w_{\sigma\gamma\tau\nu}C_{\mu\beta}\ \ ,$ (14)
with $C_{\mu\tau;\ \nu}$ being the covariant derivative of $C_{\mu\tau}$ in standard GR when there is no torsion (i.e. $a_{T}=0$, in our case).
From eq.(14), the metric compatibility can be easily proved by showing that $D_{\alpha}g_{\mu\nu}=0$. In our case, the contortion tensor $K^{\ \ \mu}_{\alpha\beta}$ and $S_{\alpha\beta\gamma}$ have the expressions,
@@
The important point to note is that the torsion vanishes if either $a_{T}=0$ or if the scalar field $\chi(x)$ that corresponds to the geometrodynamical four form is independent of both space as well as time. In the remaining portions of the article, for the sake of simplicity, we will concentrate on the case $a_{T}=0$ while assuming $\frac{\partial\chi}{\partial x^{\alpha}}$ to be non-vanishing. It is straightforward to extend the analysis to the $a_{T}\neq 0$ case in order to study the effects of torsion. For the rest of the sections, we have chosen a unit system in which $\hbar=c=1$ and the metric signature: (+ - - -).
@@
When $\chi$ satisfies the equation of motion given by eq.(21), its energy-momentum tensor (eq.(22)) gets simplified to,
@@
Now we take up the inclusion of (3+1)space-time-dimensional Chern-Simons (CS) terms in the theory. Our CS extensions of GR have been inspired by the seminar paper of Jackiw and Pi Jackiw , except for a crucial difference that they had introduced an external fixed four-vector $v_{\mu}$ instead of a dynamical $\phi_{;\mu}$. Because of this, their model had a manifest violation of Lorentz invariance Jackiw . In our case, both general as well as Lorentz covariances are maintained throughout.
First we consider a Chern-Simons (CS) term that couples electromagnetic field to the four-form $\tilde{w}$,
@@
where J is a dimensionless constant and $F_{\mu\nu}=A_{\nu,\mu}-A_{\mu,\nu}$. The action expressed in eq.(24) is invariant under diffeomorphism as well as the gauge transformations, $A_{\mu}\rightarrow A_{\mu}+\partial_{\mu}\chi$.
@@
Application of the above CS formulation to study its effects on magnetohydrodynamics as well as the $\vec{E}.\vec{B}\neq 0$ regions of a pulsar magnetosphere as a source of propagating $\tilde{w}$ have been considered pdg-ref-47 .
@@
Assuming that the geometrodynamical four-form is a dark energy candidate, the above dynamical equations were studied in an earlier work to address the observed late time acceleration of the expansion rate of the universe (pdg-ref-20 and the references therein). Our objective, in this article, is somewhat different since we propose here that the particle associated with $\tilde{w}$ is a pseudo-scalar which acts as a cold dark matter (CDM) candidate aiding in the formation of supermassive black holes when the universe was young. This is subject of Section III.
@@
Dynamical equations of motion that ensue from variations of $\mathcal{A}$ (eq.(60)) with respect to $h_{\oplus}$, $h_{\otimes}$ and $\psi$ are given by,
@@
$\partial^{\mu}\partial_{\mu}\psi+\frac{1}{2}\eta^{\mu\nu}\frac{\partial\psi}{\partial x^{\mu}}\frac{\partial\psi}{\partial x^{\nu}}-\frac{a_{2}}{4a_{1}}=h_{\oplus}\bigg{[}\frac{\partial^{2}\psi}{\partial y^{2}}-\frac{\partial^{2}\psi}{\partial z^{2}}+\frac{1}{2}\bigg{(}\frac{\partial\psi}{\partial y}\bigg{)}^{2}-\frac{1}{2}\bigg{(}\frac{\partial\psi}{\partial z}\bigg{)}^{2}\bigg{]}+2h_{\otimes}\bigg{[}\frac{\partial^{2}\psi}{\partial y\partial z}+\frac{1}{2}\frac{\partial\psi}{\partial y}\frac{\partial\psi}{\partial z}\bigg{]}$ (63)63( 63 )
where $\psi\equiv\ln\chi$.
@@
Based on the invariances of the Minkowski tensor and the flat space-time Levi-Civita tensor under proper Lorentz transformations, the present study explores the possibility of extending Einstein’s geometrical theory of gravitation by including another geometrical degree of freedom $\tilde{w}$, which is simply a generalisation of the Levi-Civita symbol, in the theory. This geometrodynamical four-form $\tilde{w}$ leads not only to a dynamical torsion it also generates Chern-Simon (CS) extensions in the 3+1-dimensional space-times.
Following closely the seminal work of Jackiw and Pi Jackiw but avoiding the use of an unphysical Lorentz vector, CS coupling between electromagnetic fields and $\tilde{w}$ comes about naturally, bringing about a modification of Einstein-Maxwell equations. Using once again the method of Jackiw , a gravitational CS term is constructed that leads to a modified Cotton tensor. However, when the dynamical equation of the four-form is used in the ensuing Einstein equation, it is shown that it is the standard Cotton tensor that affects the space-time dynamics.

### v2

@@ Abstract
By envisaging a four dimensional manifold’s space-time geometry be described by a line-element, $ds^{2}=\eta_{\mu\nu}dx^{\mu}dx^{\nu}$, Hermann Minkowski had overturned the idea of treating space and time disjointly. Furthermore, the cause and effect relation, in the realm of classical physics, attained a significant elucidation in the Minkowskian framework - infinitesimally separated events (assuming a signature (+ - - -)) for which $ds^{2}<0$ ($ds^{2}>0$) are causally disconnected (causally connected). Of course, the underlying physics involved in it is that an effect essentially follows a cause due to transfer of energy from the latter to the former, and that the rate of flow of energy is limited by the speed $c$. What is usually not stressed is that, because of this causality criteria, the imaginary numbers entered physics in a fundamental manner even before the advent of quantum mechanics, i.e. the proper distance $ds$ between any two causally disconnected events is always an imaginary number.
@@
It must be pointed out that there is another tensor in the flat space-time that is invariant under proper Lorentz transformations: the totally antisymmetric Levi-Civita tensor $\epsilon_{\mu\nu\rho\sigma}$. Since, under $x^{\mu}\rightarrow\Lambda^{\mu}_{\ \nu}x^{\nu}$,
@@
and $det(\Lambda)=1$ for proper Lorentz Transformations, it is clear that the flat space-time Levi-Civita tensor does not change under arbitrary boosts and rotations of inertial frames.
Given this special status that the Minkowski metric and the Levi-Civita tensor enjoy in the Minkowskian space-time, as well as the fact that $\eta_{\mu\nu}$ metamorphoses into a dynamical space-time metric $g_{\mu\nu}(x)$ in GR, raises a pertinent point of of elevating $\epsilon_{\mu\nu\alpha\beta}$ into a dynamical field. In the present article, we explore such a possibility by positing a geometrodynamical field, $w_{\mu\nu\rho\sigma}(x)$, totally antisymmetric in its indices, that is independent of the metric tensor and that influences space-time physics. Then, the interesting feature that $\eta_{\mu\nu}$ and $\epsilon_{\mu\nu\alpha\beta}$ both share in the flat space-time becomes relevant also in arbitrary curved space-times pdg-ref-20 .
We propose in this article that this geometrodynamical field represents a new physical degree of freedom that couples universally to all matter. In the following section, we discuss some of the differential geometric aspects of $w_{\mu\nu\rho\sigma}(x)$ that may be significant in the context of space-time physics.
@@
In a coordinate basis, this four-form field may be expressed as,
@@
Since we demand that this new dynamical field be independent of the metric tensor, instead of raising the indices of $w_{\mu\nu\rho\sigma}$ using the metric tensor, the totally antisymmetric components $w^{\mu\nu\rho\sigma}$ corresponding to the four-form $\tilde{w}$ are obtained from the associated p-vector w, so that one obtains the relation, $w^{\mu\nu\rho\sigma}\ w_{\mu\nu\rho\sigma}=-4!$ Schutz80 . This implies,
@@
$\bar{\Gamma}^{\mu}_{\alpha\beta}\equiv\Gamma^{\mu}_{\alpha\beta}-a_{T}\ g^{\mu\nu}g^{\gamma\lambda}\frac{\partial\chi}{\partial x^{\lambda}}\ w_{\nu\gamma\alpha\beta}=\Gamma^{\mu}_{\alpha\beta}-a_{T}\ g^{\mu\nu}\chi^{\ ;\gamma}\ w_{\nu\gamma\alpha\beta}\ ,$ (10)
@@
Covariant derivatives of tensor fields are obtained in the usual way,
@@
$D_{\nu}B_{\mu}=\partial_{\nu}B_{\mu}-\bar{\Gamma}^{\beta}_{\mu\nu}B_{\beta}=B_{\mu;\ \nu}+a_{T}g^{\beta\sigma}\chi^{;\gamma}w_{\sigma\gamma\mu\nu}B_{\beta}\ \ ,$ (13)
@@
$D_{\nu}C_{\mu\tau}=\partial_{\nu}C_{\mu\tau}-\bar{\Gamma}^{\beta}_{\mu\nu}C_{\beta\tau}-\bar{\Gamma}^{\beta}_{\tau\nu}C_{\mu\beta}=C_{\mu\tau;\ \nu}+a_{T}g^{\beta\sigma}\chi^{;\gamma}w_{\sigma\gamma\mu\nu}C_{\beta\tau}+a_{T}g^{\beta\sigma}\chi^{;\gamma}w_{\sigma\gamma\tau\nu}C_{\mu\beta}\ \ ,$ (14)
with $C_{\mu\tau;\ \nu}$ being the covariant derivative of $C_{\mu\tau}$ in standard GR when there is no torsion (i.e. $a_{T}=0$ case). From eq.(14), the metric compatibility can be easily proved by showing that $D_{\alpha}g_{\mu\nu}=0$. Similarly, eq.(10) entails that,
$\bar{\Gamma}^{\mu}_{\alpha\beta}A^{\alpha}A^{\beta}={\Gamma}^{\mu}_{\alpha\beta}A^{\alpha}A^{\beta}$
implying that the torsion in our case is autoparallel compatible. In our case, the contortion tensor $K^{\ \ \mu}_{\alpha\beta}$ and $S_{\alpha\beta\gamma}$ have the expressions,
@@
The important point to note is that the torsion vanishes if either $a_{T}=0$ or if the scalar field $\chi(x)$ that corresponds to the geometrodynamical four form is independent of both space as well as time. For the sake of simplicity, in the remaining portions of the article we will concentrate on the case $a_{T}=0$ while assuming $\frac{\partial\chi}{\partial x^{\alpha}}$ to be non-vanishing. It is straightforward to extend the analysis to the $a_{T}\neq 0$ case in order to study the effects of torsion. For the rest of the sections, we have chosen a unit system in which $\hbar=c=1$ and the metric signature: (+ - - -).
@@
When $\chi$ satisfies the equation of motion given by eq.(21), the above energy-momentum tensor gets simplified to,
@@
In this sub-section, we construct (3+1)-dimensional Chern-Simons (CS) terms. Our CS extensions of GR follows closely the seminal paper of Jackiw and Pi Jackiw , except for a crucial difference: they had introduced an external fixed four-vector $v_{\mu}$ instead of a dynamical $\phi_{;\mu}$ in their formulation. Because of this, their model had a manifest violation of Lorentz invariance. In our case, both general as well as Lorentz covariances are maintained throughout pdg-ref-20 .
First, we consider a Chern-Simons (CS) term that couples electromagnetic field to the four-form $\tilde{w}$,
@@
where J is a dimensionless constant and $F_{\mu\nu}=A_{\nu,\mu}-A_{\mu,\nu}$. The action expressed in eq.(24) is invariant under diffeomorphism as well as gauge transformations, $A_{\mu}\rightarrow A_{\mu}+\partial_{\mu}\xi(x)$.
@@
Application of the above CS formulation to study its effects on magnetohydrodynamics as well as the $\vec{E}.\vec{B}\neq 0$ regions of a pulsar magnetosphere as a source of propagating $\tilde{w}$ have been studied earlier pdg-ref-47 .
@@
Assuming that the geometrodynamical four-form is a dark energy candidate, the above dynamical equations were studied in an earlier work to address the observed late time acceleration of the expansion rate of the universe (pdg-ref-20 and the references therein). Our objective, in this article, is somewhat different since we propose here that the particle associated with $\tilde{w}$ is a pseudo-scalar which acts as a cold dark matter (CDM) candidate aiding in the formation of supermassive black holes when the universe was young. This is the subject of Section IV.
@@
After defining $\psi(x^{\mu})\equiv\ln{(\chi(x^{\mu}))}$, the dynamical equations of motion that ensue from variations of $\mathcal{A}$ (eq.(60)) with respect to $h_{\oplus}$, $h_{\otimes}$ and $\psi$ are given by,
@@
$\partial^{\mu}\partial_{\mu}\psi+\frac{1}{2}\eta^{\mu\nu}\frac{\partial\psi}{\partial x^{\mu}}\frac{\partial\psi}{\partial x^{\nu}}-\frac{a_{2}}{4a_{1}}=h_{\oplus}\bigg{[}\frac{\partial^{2}\psi}{\partial y^{2}}-\frac{\partial^{2}\psi}{\partial z^{2}}+\frac{1}{2}\bigg{(}\frac{\partial\psi}{\partial y}\bigg{)}^{2}-\frac{1}{2}\bigg{(}\frac{\partial\psi}{\partial z}\bigg{)}^{2}\bigg{]}+2h_{\otimes}\bigg{[}\frac{\partial^{2}\psi}{\partial y\partial z}+\frac{1}{2}\frac{\partial\psi}{\partial y}\frac{\partial\psi}{\partial z}\bigg{]}\ \ .$ (63)63( 63 )
@@
Based on the invariances of the Minkowski tensor and the flat space-time Levi-Civita tensor under proper Lorentz transformations, the present study explores the possibility of extending Einstein’s geometrical theory of gravitation by including another geometrical degree of freedom $\tilde{w}$, which is simply a generalisation of the Levi-Civita symbol, in the theory. This geometrodynamical four-form $\tilde{w}$ leads not only to a dynamical torsion it also generates Chern-Simon (CS) extensions in the 3+1-dimensional space-times. Torsion, among several other interesting implications, is also important in teleparallel gravity TelepGravity .
Pursuing closely the seminal work of Jackiw and Pi Jackiw but avoiding the use of an unphysical Lorentz vector, CS coupling between electromagnetic fields and $\tilde{w}$ comes about naturally, bringing about a modification of Einstein-Maxwell equations. Adopting once again the formulation of Jackiw , a gravitational CS term is constructed that leads to a modified Cotton tensor. However, when the dynamical equation of the four-form is used in the ensuing Einstein equation, it is shown that it is the standard Cotton tensor that affects the space-time dynamics.